Evaluate the following integrals: 3 1 32/2 7 (i) dx 6°2712 dx x²x² + 4 (4x² +9312

Evaluate the following integrals
Answer #1

: S da x²5x2 + 4 put, x=2 tand dx=2seco do 2 Seco do 4 tanto.. 2 seed secodo 4 tano Cos I do Sinzo - <scoth.coseco do + (- cox 2 2 2 W 2 x44 x 0 2 tand seth Cosecoz /2² +4 te. [ C= Integrating [ C= Integrating constant] x274 te Ans) 48 x 5x²+4 4x ✓ d3로 x². xdx 2 Nu ) 3호 2 리 3 내249 Put, 4x²+9=2 8xd 22E d2 x dx 2 Z dz 33 다 의 X 4(249)를 nola 1. 8 삊 2. x 없g ( ؟ 키 Z (SH CE- Er - [e-란그루 = [9-5bJ꾸= [ 둥 - 8 - p + spe] [2] 킥 - (-) | 솔 성 353 오 ㅜ - 3 16 cl 3 2.